OpenStudy (anonymous):

The electric force between two charged objects is 0.24 newtons. If the distance between them is increased by a factor of 4, what will be the new force? 0.015 newtons 0.03 newtons 0.06 newtons 0.12 newtons 0.36 newtons

OpenStudy (nincompoop):

what does it mean by a factor of 4

OpenStudy (anonymous):

Coulomb's Law describes the electric force between charged objects. Coulomb's Law is: \[F=\frac{1}{4\pi \epsilon}*\frac{q_1*q_2}{r^2}=k*\frac{q_1*q_2}{r^2}\] The relevant part iis that: \[(electric\ force) \propto \frac{1}{r^2}\]

OpenStudy (nincompoop):

and do you know the formula? \(E_f = k\large \frac{Q_1 Q_2}{r^2} \)

OpenStudy (nincompoop):

I put Ef instead of Fe hahaha I went dyslexic

OpenStudy (shamim):

4^2=?

OpenStudy (shamim):

4^2=16

OpenStudy (shamim):

0.24/16=?